7) Given the state of stress at point A on the circular cross section a-a of the handle assembly is oz 47.11 MPa, ryz = 48.81 MPa. Show the results on the volume element shown. 800 N 400 mm 200 mm 1500 N 20 mm1000 N Section a-a
Answer #1

This is bi-axial state of stress where \sigma _{y}=0.

Maximum principal stress= (\sigma _{y}+\sigma _{z})/2 + ((\sigma _{y}-\sigma _{z})/2 + \tau _{yz}^{2})^{0.5}

Maximu principle stress = 0+47.11/2 + ((0-47.11)/2 + 48.812 ))0.5 = 23.55+54.19=77.74 Mpa

Minimum princile stress = (\sigma _{y}+\sigma _{z})/2 - ((\sigma _{y}-\sigma _{z})/2 + \tau _{yz}^{2})^{0.5} = 23.55-54.19 = -30.64 Mpa.

Maximum shear stress = ((\sigma _{y}-\sigma _{z})/2 + \tau _{yz}^{2})^{0.5} = 54.19 Mpa.